Ouestions 19-20. 1 kg of water at 80 °C is poured into a 1 kg glass cup initially at 20 °C The specific heat of glass is cglass- 840Assume cwater 9 Cglass. kg.°C 19. What will be the equilibrium temperature? (a) 32 °C (b) 45 °C (c) 68 °C (d) 74 °C 20. How much heat will the cup receive in the process? (a) 45,360 J (b) 57,290 J (c) 63,120 J (d) 77,410.J
Answer #1

19) applying heat balance,

mwater*cwater*ΔTwater = mglass*cglass*ΔTglass

1*9cglass*(353-T)=1*cglass*(T-293)

3177-9T=T-293

3470=10T

T=347 K or T= 74 C

option d)

20) Q= m*cglass*ΔTglass=1*840*(347-293)

Q= 840*54 = 45360 J

option A is correct